OpenStudy (anonymous):

Help me with solving this system?: x-5y=2 2x+y=4

OpenStudy (anonymous):

I got to y= -2/5+x/5 and y=2x+4 But don't exactly know what to do next.

OpenStudy (king):

x=5y+2 substitute in 2nd equation 2(5y+2)+y=4 10y+4+y=4 9y=0 y=0 x=5*0+2 x=2

OpenStudy (anonymous):

Thank you! :)

OpenStudy (king):

ure welcome!